OpenStudy (anonymous):

find the feet per second 300yards -------- 1 min

OpenStudy (asevilla5):

300 yds = 900 feet 900 ft are run in 60 secs unit rate = ft / sec = 900 / 60

OpenStudy (asevilla5):

ok 1 yard = 3 feet so 300 yards = 3*300 = 900 ft 1 minute = 60 seconds unit rate = distance / time =900 / 60 feet per second '/' means divided by 900/60 = 900 --- 60 so 15

OpenStudy (anonymous):

thank you so much!